40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5874     Entry time: Fri Nov 11 13:35:19 2011
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: Feedback to ETMY 

[Kiwamu, Katrin]

20111111_freq_supp.png

 

Red and blue curves: frequency fluctuation of the beat node between PSL and YARM laser.

Green and broen curves: Actuation on ETMY.  In ALS_CONTROL.adl  ETMY filter bank 4 and 5 were switched on. Gain was 0.3

 

Nice reduction of the frequency fluctuation.

 

Y axis is in volts^2 per counts. In order to go to MHz/sqrt(Hz) you have to take the square root and then times [20Volts/(2^16)counts]*[10Hz/0.04V].

 

Started to scan the cavity, but this didn't work. Green light all out of lock. IR beam was badly aligned to cavity. Now, my time is over and I have to leave you.

Thanks, for your help and the nice time.

ELOG V3.1.3-